OpenStudy (anonymous):

What is the measure of the angle between v = <4, 0> and w = <6, 1>?

OpenStudy (amistre64):

by definition: \[|\vec u||\vec v|\cos\alpha=\vec u\cdot\vec v\]

OpenStudy (amistre64):

soo, divide and inverse cosine to determine alpha

OpenStudy (anonymous):

i still dont understand

OpenStudy (amistre64):

then you may want to review your material to determine how to do a dot product operation, and also to find the length of a vector.

OpenStudy (amistre64):

otherwise, this amounts to trig identites

OpenStudy (amistre64):

say: angle c = angle b - angle a sin(c) = sin(b-a) = sinb cosa - sina cosb

OpenStudy (anonymous):

im doing an online class to get ahead in high school for i dont have to take it my senior year so i really dont have anything to study

OpenStudy (anonymous):

i understand it this far

OpenStudy (amistre64):

but you have to have some knowledge base to work with. you understand the triggy part?

OpenStudy (anonymous):

yeah i understand that part

OpenStudy (amistre64):